OpenStudy (anonymous):

A test consists of 10 multiple choice questions, each with 5 possible answers, one of which is correct. To pass the test a student must get 60% or better on the test. If a student randomly guesses, what is the probability that the student will pass the test?

OpenStudy (anonymous):

six or more right probability you get any one right is \(\frac{1}{5}=.2\) use the binomial \[P(X=6)=\binom{10}{6}.2^6\times .8^4\] \[P(X=7)=\binom{10}{7}.2^7.8^3\] \[P(X=8)=\binom{10}{8}.2^8\times .8^2\] \[P(X=9)=\binom{10}{9}.2^8\times .8\] \[P(X=10)=.2^{10}\] compute and add

OpenStudy (anonymous):

here is the first one http://www.wolframalpha.com/input/?i=%2810+choose+6%29.2^6*.8^2

OpenStudy (anonymous):

I dont understand where the 8^? came from